My problem is:
There are 33 stairs from bottom to top in each of six stairways in the football bleachers at Waycross Stadium. If each track team member is to run four complete sets up and down each stairway, how many stairs will be covered in a workout?
My work: 33x6=198 then 198x2=206 then 206x4=824

Answer by Theo(13342)

one complete set of up and down one stairway would be 66 steps.

4 complete sets of up and down one stairway would be equal to 4 * 66 = 264 steps.

6 stairways = 264 * 6 = 1584 steps in total.

your arithmetic says:

33 * 6 = 198, then 198 * 2 = 206, then 206 * 4 = 824.

looks like it should have been:

33 * 6 = 198, then 198 * 2 = 396, then 396 * 4 = 1584.

then you would have gotten the same answer i got.

i don't know how you got 198 * 2 = 206.

the answer is 1584 as best i can determine.